OpenStudy (anonymous):

What are the simplest factors of this polynomial? x3 + 4x2 – 9x – 36 Select one of the options below as your answer: A. (x – 4)(x2 + 3) B. (x + 4)(x2 + 3) C. (x + 4)(x + 3)(x – 3) D. (x – 4)(x + 3)(x – 3)

OpenStudy (ivettef365):

split in two parts so you have x^2(x+4) -9 (x+4)

OpenStudy (ivettef365):

then you have (x^2-9)(x+4)

OpenStudy (ivettef365):

do you know how to simplify (x^2-9) = a^2-b^2 = (a+b)(a-b)

OpenStudy (anonymous):

Okagy now i understand how to go about the problem now